A boat traveled for 3 hr with a 12-mph current to reach a picnic area. The return trip against the same current took 12 hr. Find the speed of the boat in still water.
Answer by mananth(16949)

Boat speed x mph
current speed 12 mph
..
against currentspeed x- 12
withcurrent x+ 12
..
Against current 12 hours
with current x+y 3 hours
Distance is same. D=r*t
12(x-12)=3(x+12 )
12x-144 =3x+36
12x-3x =144+36
9x=180
/ 9
x= 20 mph
Boat speed = 20 mph